OpenStudy (anonymous):

how do i verify sin3xcos3x = 1/2(sin6) ????

OpenStudy (shubhamsrg):

you must be meaning sin6x there?hmm?

OpenStudy (anonymous):

sin 6x = sin 2 (3x) = 2 sin3x . cos 3x

Parth (parthkohli):

\[\iff 2\sin(3x)\cos(3x)\]You may recognize that this is \(\sin(2\cdot 3x)\)

Parth (parthkohli):

\[\sin(6x) = 2\sin(3x)\cos(3x)\]Divide both sides by \(2\).

OpenStudy (anonymous):

so 1/2(sin6x) = sin3x . cos 3x

OpenStudy (anonymous):

RHS ======== LHS

OpenStudy (anonymous):

thankyou everyone1

OpenStudy (anonymous):

as we know that sin2x=2sinx.cosx...standard formula by this we can easily say that sin2(3x)=2sin3x.cos3x........ 1/2 sin2(3x)=sin3xcos3x............. by above we can say.......... 1/2(sin6) =sin3xcos3x